# Scale-Dropout: Estimating Uncertainty in Deep Neural Networks Using Stochastic Scale

### v1

@@ I Introduction
Neural networks (NNs) have revolutionized various fields of artificial intelligence in recent years due to their exceptional ability to learn complex patterns and generate accurate predictions. NN models have produced remarkable results in a variety of tasks, including image recognition [1], natural language processing [2], and are even being deployed in safety-critical applications such as autonomous vehicles, automatic medical diagnosis [3], and automated optical inspection in industrial applications [4]. Despite the achievements of NNs, they are unable to quantify the uncertainty associated with their predictions. Estimating uncertainty is essential to understand the risks associated with prediction and to make informed decisions, especially in safety-critical domains, where incorrect predictions can have severe consequences [5].
@@
Furthermore, exact Bayesian inference is often computationally intractable, necessitating the adoption of various approximation techniques. Ensemble methods [8], and variational inference [9] are among the prominent methods. However, they have high overhead as a result of storage of multiple copies of the model or they have twice the number of parameters compared to traditional NNs. On the other hand, Monte Carlo Dropout (MC-Dropout) based Bayesian approximation [10] is particularly appealing, as they have the same number of parameters as conventional NNs. However, the number of stochastic units required for their implementation is still a concern.
Moreover, in terms of hardware architectures, BayNNs are typically implemented on von Neumann architectures where the memory and computation units are physically separated. Since BayNN applications are data intensive, data movement between the processor and memory becomes the bottleneck of the entire application, leading to the memory wall problem [11, 12]. Consequently, implementing BayNN to Computation-in-Memory (CIM) architectures with emerging resistive non-volatile memories (NVMs) [13] is an attractive option to reduce their inherent costs. In the CIM architecture, the common operation of NN, matrix-vector multiplication, is carried out inside the memory where data (NN parameters) already reside. Therefore, the memory-wall problem is alleviated, leading to a highly attractive solution to accelerate BayNN at the edge. However, implementing BayNN to the CIM architecture is not straightforward due to the deterministic nature of the architecture, the limited precision of the spintronic memories, and the design of stochastic units for Bayesian inference. Also, an approach that does not make changes to the conventional memory array but only changes to the peripheral circuitry is attractive.
In this paper, we propose the Scale Dropout, a novel regularization technique and Monte Carlo Scale Dropout (MC-Scale Dropout) based BayNN for uncertainty estimation in Binary Neural Networks (BNNs) [7]. BNNs are a model compression approach (1-bit precision weights and activations) and overcome the problem of limited precision of the spintronic memories. Furthermore, we show that, by using SD during inference, an MC-Dropout-like uncertainty estimation of the NN can be obtained. Our approach aims to provide a balanced trade-off between model uncertainty and the computational constraints of edge devices. USD employs the well-known dropout technique, dropping out the entire scale vector, thus reducing memory and computational requirements without sacrificing the quality of uncertainty estimates.
@@
Compared to existing work, our method requires only one dropout module for the model, regardless of the size of the topology. Leading to a highly scalable Bayesian approach for edge applications. The predictive performance of our method was extensively evaluated for various deep learning tasks, including classification and semantic segmentation, using different data sets and network topologies. Also, the effectiveness of our approach in estimating uncertainty is evaluated on various out-of-distribution data and metrics.
The structure of the paper is as follows. Section II describes the background on Binary Neural Networks, traditional dropout techniques, spintronic memories, CIM, and discusses related papers. In Section III, we present the proposed Scale Dropout technique, explaining its design, operation, and uncertainty estimation approach. In Section V, details of hardware implementation are discussed. In the following Section VI, our experimental results are presented at the algorithmic and hardware levels. Section VII concludes the paper.
@@ II Preliminaries
II-A Uncertainty Estimation
Uncertainty estimation in neural networks is the process by which a model provides a measure of uncertainty or confidence along with its predictions. This measure typically captures the model’s belief about the output, given the input data and the learned parameters [14]. Uncertainty estimation enables NNs to express what they do not know. Traditional neural networks, despite their impressive prediction capabilities, fail to provide these crucial uncertainty estimates. There are two distinct types of uncertainty in deep learning: Aleatoric and epistemic. Aleatoric or data uncertainty, which is irreducible even with more data, arises due to the inherent noise in the data generation process. On the other hand, uncertainty resulting from a lack of knowledge is referred to as epistemic or model uncertainty. In other words, it refers to the lack of knowledge of a model. Epistemic uncertainty can be eliminated with sufficient data from an unseen region or with knowledge about sources of uncertainty. Estimating epistemic uncertainty allows the NN to be aware of what they do not know. Therefore, this paper focused on estimating the epistemic uncertainty of a model.
II-B Conventional Dropout Methods
@@
In this paper, we take a different approach to dropout. Our approach is targeted at the scale vectors of a model. Additionally, our approach does not set the dropped value of the scale vector to zero, as done in existing works, as it prevents information flow. Instead, our approach drops the entire scale vector to one.
II-C Bayesian Neural Networks
@@
is generally computationally intractable. To overcome this challenge, various approximation methods, such as Monte Carlo dropout [10], were developed to allow feasible approximations of uncertainty for BayNNs. The approximate posterior distribution is depicted as $q(y^{*}\mid x^{*},\mathcal{D})$. By introducing dropout layers into the network during training and, more importantly, keeping them active during testing, MC-Dropout effectively simulates a sampling process from the posterior distribution of the parameters $\mathrm{\boldsymbol{\theta}}$. When the dropout layers are kept active during testing, samples can be obtained from an approximate posterior.
@@
II-D Binary NNs and Scaling
BNNs have gained popularity as a model compression technique due to their ability to reduce memory requirements by 32$\sim$ 32$\times$ 32 × compared to a full-precision NN. In BNNs, the weights and activations are binarized to $-1$ or $+1$ using the $\operatorname{sign}(.)$ function. Binarization further simplifies computationally intensive matrix-vector multiplication operations into computationally cheaper bitwise XNOR and bit counting operations, thus improving computational efficiency and leading to speed-up of 58$\sim$ 58$\times$ 58 × on CPU [20].
The scaling factor ${\boldsymbol{{\alpha}}}$ is a crucial aspect of BNN to alleviate the loss of accuracy due to binarization [20]. The scaling factor is defined as analytically calculated real-valued vectors that scale the binarized weights and activations for each layer in  [20] or a single scale vector can be learned through backpropagation similar to other parameters of the model [21]. However, some BNN algorithms, such as [22] apply the scale to the weight matrix of each layer before the XNOR operation, which is possible in a CPU or GPU implementation but may not be as feasible for a spintronics-based CIM architecture. This is because, depending on the shape of the scale vector, each neuron or channel will have a different scale factor, leading to different mapping strategy requirements for each neuron or channel. Similarly, input scaling is not feasible, as the inputs are directly converted to voltages and fed into the crossbar for computation.
@@ II-F Related Works
Previous studies have investigated hardware solutions for Bayesian and Binary Neural Networks, which are usually based on CMOS technology. In the paper [26], a novel FPGA implementation is proposed that utilizes non-linear activation functions. However, this approach may be restricted when used with larger datasets. The technique described in [27] involves a CIM implementation in which the crossbar array stores the variance parameter and stochastic resistive (RRAM) devices are used to sample the probability distribution at the input of the array. This approach requires the utilization of a single random element for each input, which is not very energy efficient. In [28], the authors take advantage of the non-idealities of RRAM devices to apply Bayesian learning. In the paper by [29], an implementation of neuron dropout was proposed that takes advantage of the stochastic and deterministic features of STT-MRAM. Unfortunately, this approach requires a random number generator (RNG) for each neuron, leading to a considerable increase in power consumption. The research in [30] showed the application of a set of resistive crossbar arrays to store probabilistic weights to execute BayNN. In [31], they proposed a suitable CIM architecture for BayNN, where multiple crossbar arrays are available and one is randomly chosen for each forward pass. In Yang et al. [32], crossbar arrays were used to construct Bayesian neural networks with the help of low-barrier MTJs, resulting in a significant decrease in energy consumption. Despite the fact that memories with low energy barriers are used, they have endurance limitations that can eventually have an impact on the precision of the CIM engine. The paper in [33] presented an alternative implementation with MRAM-based crossbar arrays that can represent mean and variance. However, this approach required considerable pre-processing to encode the mean and variance in the crossbars.
@@ III-B Scale Dropout Model Description
$\mathrm{\mathbb{a}}^{(l)}$ $=\phi(\mathbf{z}^{(l)})$ (4)
@@
$\mathbf{z}^{(l)}$ $=(\operatorname{sign}(\mathrm{\mathbb{W}}^{(l)})^{\mathcal{T}}\otimes\operatorname{sign}(\mathbf{z}^{(l-1))}+\mathrm{\mathbb{b}}^{(l)})\odot\hat{{\boldsymbol{{\alpha}}}}^{(l)}$ (7)
@@
As a result, during forward propagation, the network ignores the scale factors that correspond to the dropout mask being 0, while the scaling factor retains its original value when the randomly generated dropout mask is set to 1.
Fig. 1 shows the scale dropout concept during the train and inference time.
@@ III-E Learning with Scale-Dropout
The proposed BayNN with Scale-Dropout can be trained using stochastic gradient descent, similar to standard BNN using existing algorithms such as [7, 20]. The only difference is that for each forward pass during training, we sample a scaled network by applying Scale-Dropout. The forward and backward propagation for each iteration is performed only on this scaled network network. The gradients for each parameter are averaged over the training instances of each mini-batch. The training objective combining a Bayesian approximation and Scale-Dropout is discussed in Section IV.
@@ IV Scale-Dropout as a Bayesian Approximation
In this paper, we propose Monte Calo (MC)-Scale Dropout based Bayesian approximation that uses Scale Dropout in place of the standard dropout for Bayesian inference. Our approach expands the MC-Dropout approaches [29, 10], for BNN and better efficiency with specific learning objectives. In the following section, the learning objective and how to obtain the model uncertainty for the (MC)-Scale Dropout are discussed in detail.
@@ IV-A Learning Objective
For the proposed Monte Calo (MC)-Scale Dropout objective, we introduce a regularization function for the scales $\boldsymbol{{\boldsymbol{{\alpha}}}}$. Specifically, we design a regularization function that encourages the scale factor to be positive to preserve the sign of the computed $\mathbf{z}^{(l)}$ of a layer $l$. Also, it encourages the scale factor to be centered around one, so it scales up or down the element of $\mathcal{\mathbb{Z}}$ based on their contribution to the loss.
@@ IV-B Obtaining Model Uncertainty
Here, $x^{*}$ is the test input, $\tilde{y}$ is the stochastic prediction, and $y^{*}$ is the final prediction. We refer to this Monte Carlo estimate as the MC-Scale Dropout. In practice, this is equivalent to performing $T$ stochastic forward passes through the network and averaging the results. In the literature, this is known as model averaging [10].
@@
Here, $\mu_{Y}$ and $\sigma_{Y}$ represent predictive mean $E_{q(y^{*}\mid x^{*},\mathcal{D})}(y^{*})$, and variance $\text{Var}_{q(y^{*}\mid x^{*},\mathcal{D})}(y^{*})$ from formulas 18 and 19, respectively. For sufficiently large $T$, the confidence interval can be approximated by directly calculating the $\frac{100-k}{2}$ and $\frac{100+k}{2}$ quantile (for a $K$CI) of the predictions as
@@ V-A Modelling Spintronic-based Scale Dropout
In our design, only one spintronic-based Dropout (Spin-ScaleDrop) module is designed and implemented for the entire neural network. Thus, the proposed Spin-ScaleDrop method is reused for all layers in the CIM architecture. After the computation of a layer is performed, a new dropout mask from the dropout module is sampled for the next layer. However, due to the manufacturing and infield variation of the MTJs in the Spin-ScaleDrop, the dropout probability itself becomes a stochastic variable. We model the variation as a Gaussian distribution, the mean of the distribution $\mu$ represents the expected dropout probability, and $\sigma$ represents the device variations. Therefore, the probability of dropout $p$ of a layer $l$ can be modelled as

### v2

@@ I Introduction
In recent years, Neural networks (NNs) have revolutionized various fields of artificial intelligence due to their exceptional ability to learn complex patterns and generate accurate predictions. NN models have produced remarkable results in a variety of tasks, including image recognition [1], natural language processing [2], and are even being deployed in safety-critical applications such as autonomous vehicles, automatic medical diagnosis [3], and automated optical inspection in industrial applications [4]. Despite the achievements of NNs, they are unable to quantify the uncertainty associated with their predictions. Estimating uncertainty is essential to understand the risks associated with prediction and to make informed decisions, especially in safety-critical domains, where incorrect predictions can have severe consequences [5].
@@
Furthermore, exact Bayesian inference is often computationally intractable, necessitating the adoption of various approximation techniques. The ensemble methods [8], and variational inference [9] are among the prominent methods. However, they have high overhead as a result of the storage of multiple copies of the model, or they have twice the number of parameters compared to traditional NNs. On the other hand, Monte Carlo Dropout (MC-Dropout) based Bayesian approximation [10] is particularly appealing, as they have the same number of parameters as conventional NNs. However, the number of stochastic units required for their implementation is still a concern.
Moreover, in terms of hardware architectures, BayNNs are typically implemented on von Neumann architectures, where the memory and computation units are physically separated. Since BayNN applications are data intensive, data movement between the processor and memory becomes the bottleneck of the entire application, leading to the memory wall problem [11, 12]. Consequently, implementing BayNN to Computation-in-Memory (CIM) architectures with emerging resistive non-volatile memories (NVMs) [13] is an attractive option to reduce their inherent costs. In the CIM architecture, the common operation of NN, matrix-vector multiplication, is carried out inside the memory where data (NN parameters) already reside. Therefore, the memory-wall problem is alleviated, leading to a highly attractive solution to accelerate BayNN at the edge. However, implementing BayNN to the CIM architecture is not straightforward due to the deterministic nature of the architecture, the limited precision of the spintronic memories, and the design of stochastic units for Bayesian inference. Also, an approach that does not make changes to the conventional memory array but only changes to the peripheral circuitry is attractive.
Therefore, in the literature, several methods [14, 15, 16, 17, 18, 19, 20] have been proposed to implement Dropout-based as well as other approximate BayNNs into CIM architectures with different NVM technologies. However, they have limitations, including a) requiring numerous Dropout modules (typically one for each neuron), b) not suitable for convolutional layers, and c) having high power consumption.
In this paper, we propose the Scale Dropout, a novel regularization technique and Monte Carlo Scale Dropout (MC-Scale Dropout) based BayNN for uncertainty estimation in Binary Neural Networks (BNNs) [7]. BNNs utilize 1-bit precision weights and activations, effectively addressing the constraints posed by limited-precision spintronic memories. Moreover, by integrating Scale Dropout during inference, we achieve a robust uncertainty estimation, comparable to MC-Dropout. Our approach aims to provide a balanced trade-off between model uncertainty and the computational constraints of edge devices. The scale Dropout employs the proposed vector-wise unitary Dropout technique, dropping out the entire scale vector, thus reducing memory and computational requirements without sacrificing the quality of uncertainty estimates. The scale vector, an additional parameter in BNN, is crucial in modern BNN algorithms to reduce quantization error.
@@
Compared to existing work, our method requires only one Dropout module for the model, regardless of the size of the topology leading to a highly scalable Bayesian approach for edge applications. The predictive performance of our method was extensively evaluated for various deep learning tasks, including classification and semantic segmentation, using different data sets and network topologies. Also, the effectiveness of our approach in estimating uncertainty is evaluated on various out-of-distribution data and metrics.
The structure of the paper is as follows: Section II describes the background on Binary Neural Networks, traditional Dropout techniques, spintronic memories, CIM, and discusses related papers. In Section III, we present the proposed Scale Dropout technique, explaining its design, operation, and uncertainty estimation approach. In Section V, details of hardware implementation are discussed. In Section VI, simulation and experimental setup are discussed, and in Section VII, our experimental results are presented at the algorithmic and hardware levels. Later, in Section VIII, we discuss various aspects of the proposed method, and Section IX concludes the paper.
@@ II Preliminaries
II-A Binary NNs and Scaling
BNNs have gained popularity as a model compression technique due to their ability to reduce memory requirements by 32$\sim$ 32$\times$ 32 × compared to a full-precision NN. In BNNs, the weights and activations are binarized to $-1$ or $+1$ using the $\operatorname{sign}(.)$ function. Binarization further simplifies computationally intensive matrix-vector multiplication operations into computationally cheaper bitwise XNOR and bit counting operations, thus improving computational efficiency and leading to a speed-up of 58$\sim$ 58$\times$ 58 × on CPU [21].
The scale vector ${\boldsymbol{{\alpha}}}$ is a crucial aspect of BNN to alleviate the loss of accuracy due to binarization [21]. Here, a real-valued vector multiplies the weighted sum, weights, or activations of a layer. The scale vector can be defined in two ways, such as analytically calculated values that scale the binarized weights and activations for each layer in [21] or it can be learned through backpropagation similar to other parameters of the model [22]. In terms of the location of the scale vector, applying the scale to the weight matrix of each layer before the XNOR operation (as done in [23]) is possible in a CPU or GPU implementation but may not be as feasible for a spintronics-based CIM architecture. This is because, depending on the shape of the scale vector, each neuron or channel will have a different scale factor, leading to different mapping strategy requirements for each neuron or channel. Similarly, input scaling is not feasible, as the inputs are directly converted to voltages and fed into the crossbar for computation.
II-B Uncertainty Estimation
Uncertainty estimation in neural networks is the process by which a model provides a measure of uncertainty or confidence along with its predictions. This measure typically captures the model’s belief about the output, given the input data and the learned parameters [24]. Uncertainty estimation enables NNs to express what they do not know. Traditional neural networks, despite their impressive prediction capabilities, fail to provide these crucial uncertainty estimates. There are two distinct types of uncertainty in deep learning: Aleatoric and epistemic [25]. Aleatoric or data uncertainty, which is irreducible even with more data, arises due to the inherent noise in the data generation process. On the other hand, uncertainty resulting from a lack of knowledge is referred to as epistemic or model uncertainty. In other words, it refers to the lack of knowledge of a model. Epistemic uncertainty can be eliminated with sufficient data from an unseen region or with knowledge about sources of uncertainty. Estimating epistemic uncertainty allows the NN to be aware of what they do not know. Therefore, this paper focused on estimating the epistemic uncertainty of a model.
II-C Conventional Dropout Methods
@@
In this paper, we take a different approach to Dropout. Our approach is focused on the scale vectors of a binary NN model. Additionally, our approach does not set the dropped value of the scale vector to zero, as done in existing work, as it prevents information flow. This means that the whole weighted sum of a layer would be zeroed out if the scaled vector were set to zero. Instead, our approach drops the entire scale vector to one. That means in the dropped behavior, the scale vector is bypassed with a probability, and the weighted sum value of a layer remains unchanged.
II-D Bayesian Neural Networks
@@
is generally computationally intractable. To overcome this challenge, various approximation methods, such as Monte Carlo Dropout [10], were developed to allow feasible approximations of uncertainty for BayNNs. The approximate posterior distribution is depicted as $q(y^{*}\mid x^{*},\mathcal{D})$.
By introducing Dropout layers into the network during training and, more importantly, keeping them active during testing, MC-Dropout effectively simulates a sampling process from the posterior distribution of the parameters $\mathrm{\boldsymbol{\theta}}$. When the Dropout layers are kept active during testing, samples can be obtained from an approximate posterior.
@@ II-F Related Works
Previous studies have investigated hardware solutions for Bayesian and Binary Neural Networks, which are usually based on CMOS technology. In the paper [34], a novel FPGA implementation is proposed that utilizes non-linear activation functions. However, this approach may be restricted when used with larger datasets. With similar technology, work in [35] proposed an architecture to implement MC-Dropout. The study described in [36] introduces an approach where only a partial BayNN is implemented, treating solely the last layers of the network as Bayesian. The technique described in [37] involves a CIM implementation in which the crossbar array stores the variance parameter and stochastic resistive RRAM devices are used to sample the probability distribution at the input of the array. This approach requires a single random element for each input, which is not very energy efficient. In [38], the authors take advantage of the non-idealities of RRAM devices to apply Bayesian learning. In the paper by [17], an implementation of neuron Dropout was proposed that takes advantage of the stochastic and deterministic features of STT-MRAM. Unfortunately, this approach requires a random number generator (RNG) for each neuron, leading to a considerable increase in power consumption. The research in [18] showed the application of a set of resistive crossbar arrays to store probabilistic weights to execute BayNN. In [14], they proposed a suitable CIM architecture for BayNN, where multiple crossbar arrays are available and one is randomly chosen for each forward pass. In Yang et al. [39], crossbar arrays were used to construct Bayesian neural networks with the help of low-barrier MTJs, resulting in a significant decrease in energy consumption. Despite the fact that memories with low-energy barriers are used, they have endurance limitations that can eventually have an impact on the precision of the CIM engine. The paper in [40] presented an alternative implementation with MRAM-based crossbar arrays that can represent mean and variance. However, this approach required considerable pre-processing to encode the mean and variance in the crossbars.
@@ III-B Scale Dropout Model Description
$\mathrm{\mathbb{a}}^{(l)}$ $=\phi(\hat{\mathbf{z}}^{(l)})$ (4)
@@
$\mathbf{z}^{(l)}$ $=(\operatorname{sign}(\mathrm{\mathbb{W}}^{(l)})^{\top}\otimes\operatorname{sign}(\mathbf{z}^{(l-1)})+\mathrm{\mathbb{b}}^{(l)})\odot\hat{{\boldsymbol{{\alpha}}}}^{(l)}$ (7)
@@
As a result, during forward propagation, the network ignores the scale factors that correspond to the Dropout mask being 0, while the scaling factor retains its original value when the randomly generated Dropout mask is set to 1. Fig. 1 shows the scale Dropout concept during the train and inference time.
@@ III-E Learning with Scale-Dropout
The proposed BayNN with Scale-Dropout can be trained using stochastic gradient descent, similar to standard BNN using existing algorithms such as [7, 21]. The only difference is that for each forward pass during training, we sample a scaled network by applying Scale-Dropout. The forward and backward propagation for each iteration is performed only on this scaled network. The gradients for each parameter are averaged over the training instances of each mini-batch. The training objective combining a Bayesian approximation and Scale-Dropout is discussed in Section IV.
@@ IV Scale-Dropout as a Bayesian Approximation
In this paper, we propose Monte Calo (MC)-Scale Dropout based Bayesian approximation that uses Scale-Dropout in place of the standard Dropout for Bayesian inference. Our approach expands the MC-Dropout approaches [17, 10], for BNN and better efficiency with specific learning objectives. In the following section, the learning objective and how to obtain the model uncertainty for the (MC)-Scale Dropout are discussed in detail.
@@ IV-A Learning Objective
For the proposed Monte Calo (MC)-Scale Dropout objective, we introduce a regularization function for the scales $\boldsymbol{{\boldsymbol{{\alpha}}}}$. Specifically, we design a regularization function that encourages the scale factor to be positive to preserve the sign of the computed $\mathbf{z}^{(l)}$ of a layer $l$. Also, it encourages the scale factor to be centered around one, so it scales up or down the element of $\mathbf{z}$ based on their contribution to the loss.
@@ IV-B Obtaining Model Uncertainty
Here, $x^{*}$ is the test input, $q(y^{*}\mid x^{*},\mathcal{D})$ is the posterior distribution, $\tilde{y}$ is the stochastic prediction, and $y^{*}$ is the final prediction. We refer to this Monte Carlo estimate as the MC-Scale Dropout. In practice, this is equivalent to performing $T$ stochastic forward passes through the network and averaging the results. In the literature, this is known as model averaging [10].
In terms of the posterior distribution of the output, equations 5 can be modified as
$\mathbf{z}^{(l)}$ $=S^{(l)}\odot\text{diag}(d)$ (19)
$d^{(l)}$ $\sim\text{Bernoulli}(p)\text{ for }l=1,\cdots,L$ (20)
Here, $S$ represents the weighted sum of a layer. Batch normalization is applied to $\mathbf{z}^{(l)}$. Thus, the sampling process of the Dropout mask is the same as that of the MC-Dropout. In an empirical evaluation (shown later in the VIII), we observed that the distribution of output for each class approaches a Gaussian distribution as the number of stochastic forward passes (T) through the network increases. This is due to the aggregate effect of the scalar dropout mask over many forward passes, which can be seen as introducing a form of multiplicative noise.
@@
Here, $\mu_{Y}$ and $\sigma_{Y}$ represent predictive mean $E_{q(y^{*}\mid x^{*},\mathcal{D})}(y^{*})$, and variance $\text{Var}_{q(y^{*}\mid x^{*},\mathcal{D})}(y^{*})$ from formulas 18 and 21, respectively. For sufficiently large $T$, the confidence interval can be approximated by directly calculating the $\frac{100-k}{2}$ and $\frac{100+k}{2}$ quantile (for a $K$CI) of the predictions as
@@ V-A Modelling Spintronic-based Scale Dropout
In our design, only one spintronic-based Dropout (namd here Spin-ScaleDrop) module is designed and implemented for the entire neural network. Thus, the proposed Spin-ScaleDrop module is reused for all layers of the CIM architecture. After the computation of a layer is performed, a new Dropout mask from the Spin-ScaleDrop Module module is sampled for the next layer. However, due to the manufacturing and infield variation of the MTJs in the Spin-ScaleDrop, the Dropout probability itself becomes a stochastic variable. We model the variation as a Gaussian distribution, the mean of the distribution $\mu$ represents the expected Dropout probability, and $\sigma$ represents the device variations. Therefore, the probability of Dropout $p$ of a layer $l$ can be modelled as